Mitts Cosmetics Co.'s stock price is $52.65, and it recently paid a $1.50 dividend. This dividend is expected to grow by 24% for the next 3 years, then grow forever at a constant rate, g; and rs = 11%. At what constant rate is the stock expected to grow after Year 3? Round your answer to two decimal places. %
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

в C $52.65 Current Price of stock (PO) required return (rs) Last Dividend(DO) Growth rate for next 3 years 11% $1.50 24% Year

Cell reference -

B Current Price of stock (PO) required return (rs) Last Dividend(DO) Growth rate for next 3 years 52.65 0.11 1.5 0.24 Year Di

Please note:

D3*(1 +9) Terminal value = r-9

64.85 = 2.31 *(1+9) 0.11 - 9

g= 64.85 * 0.11 - 2.31 = 64.85 + 2.31

g= 7.19%

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.